Solve the differential equation by variation of parameters, subject to the initial conditions
y(0) = 1, y'(0) = 0.

9y'' − y = x(e^x/3)

y(x) = ____
1
Expert's answer
2020-06-29T18:52:38-0400

The auxiliary equation is "9m^2+1=0", which gives "m = \pm\dfrac{1}{3}."

The complementary function is "C.F = c_{1}e^{-\frac{x}{3}}+c_{2}e^{\frac{x}{3}}" .

The particular integral is given by,

"P.I = Pf_{1}+Qf_{2}" , where "f_{1} = e^{-\frac{x}{3}}, f_{2} = e^{\frac{x}{3}}" and


"P=-\displaystyle\int \dfrac{f_{2}X}{W(f_{1},f_{2})}dx", "Q=\displaystyle\int \dfrac{f_{1}X}{W(f_{1},f_{2})}dx~ \text{and}~ X = xe^{\frac{x}{3}}".

Now,

"W(f_{1},f_{2}) = f_{1}f'_{2}-f'_{1}f_{2} \\ ~~~~~~~~~~~~~~~~~~= e^{-\frac{x}{3}} \cdot \frac{e^{\frac{x}{3}}}{3} - \frac{-e^{-\frac{x}{3}}}{3} \cdot e^{\frac{x}{3}} = \frac{2}{3}"


"P=-\displaystyle\int \dfrac{f_{2}X}{W(f_{1},f_{2})}dx
=-\displaystyle\int \dfrac{e^{\frac{x}{3}} \cdot xe^{\frac{x}{3}}}{\frac{2}{3}}dx\\
~~~~=-\dfrac{3}{2}\displaystyle\int xe^{\frac{2x}{3}}dx =-\dfrac{3}{2}\biggl\{\frac{xe^{\frac{2x}{3}}}{\frac{2}{3}} - \dfrac{e^{\frac{2x}{3}}}{\frac{4}{9}}\biggr\}\\
~~~~=\dfrac{3}{2}\biggl\{\dfrac{9}{4}{e^{\frac{2x}{3}}} - \dfrac{3}{2}{xe^{\frac{2x}{3}}} \biggr\}=\dfrac{9}{8}e^{\frac{2x}{3}}(3-2x)"


"Q=\displaystyle\int \dfrac{f_{1}X}{W(f_{1},f_{2})}dx = \displaystyle\int \dfrac{e^{-\frac{x}{3}} \cdot xe^{\frac{x}{3}}}{\frac{2}{3}}dx = \dfrac{3}{4}x^{2}"


Therefore,

"P.I = Pf_{1}+Qf_{2} \\~~~~~~~= \dfrac{9}{8}e^{\frac{2x}{3}}(3-2x) \cdot e^{-\frac{x}{3}} + \dfrac{3}{4}x^{2} \cdot e^{\frac{x}{3}} \\~~~~~~~~~=\dfrac{3}{8}e^{\frac{x}{3}}(2x^{2}-6x+9)"

Thus,

"y = c_{1}e^{-\frac{x}{3}} + c_{2}e^{\frac{x}{3}} + \dfrac{3}{8}e^{\frac{x}{3}}(2x^{2}-6x+9)"


Given the initial condition, "y(0) = 1, y'(0)=0."

Using these conditions we get,

"c_{1}+c_{2} = -\frac{19}{8}\\~
-c_{1}+c_{2}=\frac{27}{8}"

Solving, we get "c_{1} = -\frac{23}{8}, c_{2}=\frac{1}{2}" and hence


"y = -\dfrac{23}{8}e^{-\frac{x}{3}} + \dfrac{1}{2}e^{\frac{x}{3}} + \dfrac{3}{8}e^{\frac{x}{3}}(2x^{2}-6x+9)\\
~\\~~~=\dfrac{1}{8}e^{\frac{x}{3}}(6x^{2}-18x+31)-\dfrac{23}{8}e^{-\frac{x}{3}}"